Sol Summary – November 28th

Sol: 1
Summary Title: Training & Experiments Initialization
Author’s name: Matt Eby
Mission Status: Training
Sol Activity Summary: Hab/rover training, deployment of mirror experiment, EVA tool-demo, setup of exercise protocol equipment, and project phantom initialization.
Look Ahead Plan: Rerun one Covid test, continue with project phantom checkout, complete training, and first suited EVA tomorrow.
Anomalies in work: N/A
Weather: sunny/high level clouds in the morning, followed by elevated winds mid-day and increasing clouds by dusk.
Crew Physical Status: Crew is adjusting to our new environment, have healthy appetites and are generally in good spirits.
Reports to be filed: Journalist, Commander, Operations, Photos, EVA request (No Green Hab report since crew did not access, HSO checklist pending access to remaining modules)
Support Requested: N/A

Operations Report – November 28th

SOL: 1

Name of person filing report: Ashley Kowalski

Non-nominal systems: The red webbing strap that is used to secure the front airlock door broke in half. Also, the two silver knobs on the oven do not work well. Only the black knobs work consistently.

Notes on non-nominal systems: Crew temporarily repaired broken red webbing strap with a water knot, and it is currently holding. As for the oven knobs, the crew switches them as needed.

Spirit rover used: (Used for rover training today)
Hours: 201.5 (start), 201.7 (end)
Beginning charge: 100%
Ending charge: 90%
Currently charging: No

Opportunity rover used: (Used for rover training today)
Hours: 105.4 (start), 105.5 (end)
Beginning charge: 100%
Ending charge: 86%
Currently charging: No

Curiosity rover used: (Used for rover training today)
Hours: 214.4 (end) (starting number was not noted)
Beginning charge: Not noted
Ending charge: 89%
Currently charging: No

Perseverance rover used: Not used
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A
General notes on rovers: No issues noted with any of the rovers used today.

Summary of Hab operations:
WATER USE: ~25 gallons
Water (static tank): ~525 gallons
Static tank pipe heater (on or off): On
Static tank heater (On or off): On
Toilet tank emptied: No
Summary of internet: No issues with the internet. Connection is good.
Summary of suits and radios: Two of our radios do not work once they are taken off the charging ports.
Summary of GreenHab operations: GreenHab is not yet accessible to the crew.
WATER USE: Unknown
Heater: Unknown
Supplemental light: Unknown
Harvest: Unknown
Summary of ScienceDome operations: ScienceDome is not yet accessible to the crew.
Dual split: N/A
Summary of RAM operations: No work was done in the RAM today.
Summary of any observatory issues: No observatory issues.
Summary of health and safety issues: No health and safety issues.

Questions, concerns and requests to Mission Support: The password for the Google drive account to upload videos does not work. We are trying the password that was provided in the most recent handbook. Please let the crew know the password when you have a moment. Thanks!

Journalist Report – November 28th

Author’s name: Dr. Kristine Ferrone, Commander

Sol 1 dawned bright and sunny on Mars. Crew 269 awoke one by one and enjoyed breakfasts of hot oatmeal with cinnamon apples and homemade bread with honey. The crew commander, Dr. Kristine Ferrone, owns a mobile coffee shop back on Earth and prepared freshly percolated espresso lattes for the crew with oat milk and agave syrup. After breakfast, the health and safety officer, Barbara Braun, measured each crew member’s temperature, pulse, and blood pressure and privately discussed each crew member’s mental and physical health as they adapt to their new environment. The crew is generally in good health and good spirits, adjusting well to life on Mars, partially due to a special communications protocol via the Discord app (programmed by Barbara’s teenage son, Elias) that allows them to chat with their friends, family, colleagues, and education/outreach contacts throughout the mission with a built-in Mars to Earth time delay. After morning activities, the crew then split into teams to prepare their experiments for operations throughout the mission. Kristine and the executive officer, Alli Taylor, explored the area around the Habitat with tools they brought from Earth for collecting samples of Martian regolith and taking measurements of the ambient radiation environment. The crew engineer, Ashley Kowalski, deployed a set of mirror coating samples to a particularly windy, dusty, and sunny location near the Habitat. These samples will be exposed to the harsh environment then collected and analyzed under a microscope on a later Sol with results sent back to the Primary Investigator on Earth, Dr. Chelsea Appleget. The technology officer, Trevor Jahn, performed initial setup and calibration of his Project Phantom virtual/augmented reality hardware for experiments on future Sols. Ashley, Barbara, and green hab officer, Matt Eby, trained on the use of the DSLR camera provided by The Aerospace Corporation and practiced taking photos around the Habitat. Barbara and Matt also set up and tested a compact fitness kit for the crew to use for exercise throughout the mission. Barbara is an experienced breadmaker back on Earth and, in the afternoon, she began the preparation of sourdough bread for the crew to enjoy with future meals. In the later afternoon, the crew learned all about living in the Habitat and operating its systems from the station director, Dr. Shannon Rupert. Topics such as how to operate the Martian toilet, shower, and kitchen were covered so that the crew can live comfortably in the station. Upon completion of their Habitat training, the crew met the assistant director, Sergii Iakymov, for training on the rovers. The crew had a great time learning how to operate and drive the rovers around the Martian surface and drove to the Marble Ritual area near the Habitat. Each crew member had an opportunity to drive the rover and learn how to navigate the Martian terrain. Upon completion of their training, the crew returned to the Habitat to write their reports and prepare the evening meal. Alli and Barbara prepared a meal of chili with crumbled beef and homemade tortillas.

Commander Report – November 28th

Sol: 1
Summary Title: Training & Initial Experiments
Author’s name: Dr. Kristine Ferrone, Commander
Mission Status: Nominal
Commander’s Report:

Crew 269 awoke this morning to another day of beautiful weather on Mars, sunny and clear. Winds picked up later in the day but remained manageable for work outside the Hab. The crew enjoyed oatmeal, homemade bread, and espresso for breakfast, then HSO recorded crew health statistics. The crew then split up to perform pre-sim experiment preparations. CDR and XO demonstrated the use of EVA tools for regolith sample collection and radiation environment measurements. Engineer deployed a set of mirror coating samples to be exposed to the light and dust environment to be collected on EVA and analyzed on a later Sol. Technology Officer performed set up operations for Project Phantom VR/AR experiment to be conducted on EVAs on future Sols. Green Hab Officer and HSO set up and tested a compact fitness kit for use on future Sols for elective crew exercise. Green Hab Officer, HSO, and Engineer trained on the use of DSLR camera and practiced taking photos around the Hab. In the afternoon, the crew completed Hab operations and rover training with Shannon and Sergii. The crew finished the evening with a dinner of chili and homemade tortillas.

Copyright © The Mars Society. All rights reserved. | Main Site